# Over-the-Air Federated Adaptive Data Analysis: Preserving Accuracy via Opportunistic Differential Privacy

### v1

@@ Abstract
Adaptive data analysis (ADA) involves a dynamic interaction between an analyst and a dataset owner, where the analyst submits queries sequentially, adapting them based on previous answers. This process can become adversarial, as the analyst may attempt to overfit by targeting non-generalizable patterns in the data. To counteract this, the dataset owner introduces randomization techniques, such as adding noise to the responses. This noise not only helps prevent overfitting but also enhances data privacy. However, it must be carefully calibrated to ensure that the statistical reliability of the responses is not compromised. In this paper, we extend the ADA problem to the context of distributed datasets. Specifically, we consider a scenario where a potentially adversarial analyst interacts with multiple distributed responders through adaptive queries. We assume that the responses are subject to noise introduced by the channel connecting the responders and the analyst. We demonstrate how, through a federated mechanism, this noise can be opportunistically leveraged to enhance the generalizability of ADA, thereby increasing the number of query-response interactions between the analyst and the responders. We illustrate that careful tuning of the transmission amplitude, based on the theoretically achievable bounds, can significantly impact the number of accurately answerable queries.
@@ I Introduction
In the era of big data, the necessity for systems that can dynamically and efficiently analyze vast and complex datasets has never been greater. One prominent approach in this domain is Adaptive Data Analysis (ADA), where the analyst aims to derive insights about an unknown distribution over a given domain by adaptively exchanging a series of query-answer pairs with a responding mechanism that possesses a dataset of samples drawn from that distribution. However, theoretical considerations of ADA in contexts involving distributed datasets is relatively unexplored, despite the growing prevalence of such datasets in modern computing environments. This paper suggests that implementing ADA within distributed data centers can be seen as an extension of Federated Learning, which traditionally focuses on using gradient descent queries to train deep neural networks across decentralized locations. The following subsections offer a necessary examination of ADA and Federated Learning.
@@ I-A Adaptive Data Analysis
In the basic model of ADA, there is one analyst ($\mathcal{A}$) and one answering mechanism ($\mathcal{M}$), communicating with each other through error-free channels with unlimited capacities. Hence, there is no concern about the practical communication limitations in this basic form of the problem. It’s assumed that the answering mechanism $\mathcal{M}$ has access to the dataset $S$ with $n$ samples generated from the domain $\mathcal{X}$ by the distribution $\mathcal{D}_{\mathcal{X}^{n}}$. Albeit, in the general case, $\mathcal{D}_{\mathcal{X}^{n}}$ is not necessarily a product distribution but in the sequel we add this assumption, as it’s the case in most of the basic papers in the field. So we assume that $\mathcal{D}_{\mathcal{X}^{n}}\equiv\mathcal{P}^{n}$, where $\mathcal{P}$ is a distribution on $\mathcal{X}$. The analyst $\mathcal{A}$ doesn’t have direct access to $S$ but can ask queries about it from $\mathcal{M}$. Various families of queries are considered in the ADA papers, among which the most basic one is the statistical queries, studied in the pioneer work [1] and also in several subsequent works such as [2, 3]. In this paper we work with this family of queries, too. The statistical queries are defined in the below.
@@
Despite their simplicity, the statistical queries cover a substantial range of practical queries. For example, gradient descent updates are indeed a vector of statistical queries. Note that the most of federated learning papers take their attention on the gradient descent queries, because of their pivotal role in the learning procedure of deep neural networks. Thus, from another perspective, our work can be seen as an abstraction of those studies.
The analyst $\mathcal{A}$ can make his queries in an adaptive manner based on the previous question-answers. More exact, the $i$’th query, i.e. $q_{i}$, is asked considering the sequence $(q_{1},a_{1},\dots,q_{i-1},a_{i-1})$, in which $a_{j}$ is $\mathcal{M}$’s answer to the $j$’th question, using the dataset $S$. It’s desired to keep the empirical answers close to the accurate answers based on the underlying distribution. The following definition, gives the exact formulation of the considered accuracy criterion.
@@
In the basic model, because $\mathcal{M}$ does not know the action model of $\mathcal{A}$, it considers the worst-case scenario. Specifically, $\mathcal{M}$ assumes that $\mathcal{A}$ selects queries adversarially, aiming to overfit to the dataset $S$ and thereby violate the accuracy criterion defined by (2). On the other hand, $\mathcal{M}$ aims to answer as much query as possible provided that (2) stays valid. So he looks for an efficient answering mechanism to achieve this goal. If the queries were non-adaptive, i.e. chosen independent of observations from data, a natural unbiased action by $\mathcal{M}$ would be to respond with the empirical answers calculated on $S$. But there is well-known examples which shows this can be an inefficient choice for the adaptive situation. For example see [1]. A widely adopted strategy in the research community for handling the adaptive case is to randomize empirical answers by introducing stochastic noise. The concept of adding noise originates from the data privacy framework, particularly differential privacy, which functions as a privacy-enhancing mechanism [4]. In this solution, i.i.d. noise samples from an appropriate distribution are generated and added to each empirically calculated answer. The noise distribution is typically chosen to be either Gaussian or Laplace. In this paper, we proceed with the Gaussian distribution, as we aim to leverage Gaussian channel noise as the source of randomization. The variance of the randomization noise is calibrated to ensure that the accuracy criterion in (2) is satisfied.
Since 2015, numerous papers have explored various aspects of Adaptive Data Analysis (ADA). In this study, we will not delve into the specifics of their theoretical findings; instead, our focus is on applying one of these foundational results to our proposed approach. To the best of our knowledge, no prior work has addressed the physical channel aspects of this issue. Furthermore, this study is the first to investigate the federated scenario within the context of ADA.
@@ I-B Federated learning
As mentioned earlier, previous studies have not examined the physical and distributed aspects of ADA. However, the concept of distributed responders in our work is closely related to federated learning (FL). Accordingly, we reference selected works from the extensive FL literature that address aspects similar to those discussed in our research.
In Federated Learning (FL), a central node, referred to as the Parameter Server (PS), communicates with multiple edge nodes known as clients. All nodes collaborate toward the shared goal of learning the weights of a neural network. Each client has access to a local dataset and contributes to the training process using this data. At the beginning of each training round, the PS broadcasts the latest set of weights to all nodes. Each node then computes the gradient of the loss function based on the current parameter values and its local dataset. Finally, the nodes send their local gradient estimations to the PS, which aggregates them to produce a global, more accurate estimation.
In the literature, various channel models and communication methods have been proposed for the gradient transmission step. These communication approaches are generally categorized into two groups: analog transmission and digital transmission. In this paper, we adopt the analog approach. In analog transmission, clients simultaneously send their estimates to the PS using uncoded transmission. This method is bandwidth-efficient due to the shared use of the channel. Fortunately, due to the additive property of the medium, local estimations naturally aggregate on the air. While this offers computational and storage efficiency, the aggregated signal may carry more noise and interference compared to digital transmission. However, these effects are not always detrimental. Some federated learning studies have leveraged channel distortions to enhance client data privacy. Moreover, [5] shows that although channel imperfections may slow the convergence rate, they can improve the accuracy of the trained network. Thus, there appears to be a dual relationship between data privacy and algorithm accuracy.
Another crucial aspect in FL context is the channel model. One key factor is the Channel State Information (CSI). In [6, 7, 8, 9], perfect global CSI for the PS and perfect local CSI for the edge devices are assumed. In [10, 11, 12, 13, 14, 15], perfect local CSI at the edge devices is assumed. In [16], each edge device is assumed to have imperfect knowledge of its channel gain to the PS, meaning there is some mismatch between its estimation and the exact channel gain. Another type of imperfect CSI is explored in [15, 17], where the PS is assumed to know only the sum of the fading coefficients, and the edge devices have no CSI knowledge. In the scheme proposed by [5], no CSI is required at all.
@@
For a more practical approach, several works, including [13, 10, 12], consider fading effects, specifically the block-flat fading regime, where the channel condition remains constant during each communication round and may change afterward. In [6], fast fading is considered. Also [11] incorporates both path loss and small-scale fading into the model and assumes that the EP’s have perfect knowledge about both the losses.
@@ I-C Contributions and Paper Structure
In this study, we consider, for the first time, performing ADA through noisy communication channels. As previously mentioned, in the ADA framework, independent noise samples are typically added to answers to maintain statistical validity. Therefore, rather than viewing channel noise as a limitation, we can opportunistically leverage it as a substitute for artificial noise.
Furthermore, we extend this approach to a distributed scenario, where an analyst—potentially adversarial—interacts with multiple EPs via a sequence of queries and responses, rather than working with a single answerer. The responses are transmitted through an Additive White Gaussian Noise (AWGN) channel. We propose a response mechanism that exploits the channel noise to enhance the generalizability of ADA and optimize the number of query-response interactions.
@@ II System Model
As depicted in Fig. 1, a distributed situation of ADA is considered. In this problem we assume that each of the $L$ edge points $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\dots$, and $\mathcal{M}_{L}$ are connected to a central point named $\mathcal{A}$ via a communication channel with a limited capacity. Each edge point (EP), e.g. $\mathcal{M}_{l}$, holds a local dataset $S_{l}$ including $n_{l}$ data samples and can act as an answering mechanism. The central point (CP) broadcasts its queries via error-free channels to all or some of them and receives the answers from them through a wireless MAC. Note that, this channel model is similar to the one proposed in lots of recent federated learning works, among which we refer to some of them in the previous section. This scenario can model a network of sensors communicating with a central point, which aim to get some inferences about the environment via data samples collected by the sensors.
More exact, in our considered scheme, the answering EPs transmit an analog signal with amplitude proportional to the answer of query and let the noisy channel to play the role of the stabilizing mechanism. Due to the independence of the noise samples, in each round of the transmission, the answers experience a new randomness, making the situation consistent with the popular solution for controlling the bias in the basic model of adaptive data analysis problem. This idea is also applied in some recent federated learning works to preserve the data privacy of the EPs.
@@ III Point-to-Point ADA without any communication constraint
In this section we begin with quoting a key theorem about the basic model of the ADA from [3]. This theorem relates the variance of the Gaussian noise to the accuracy parameters of the answering mechanism. Note that our representation of the theorem is a bit different from the reference, but equivalent to it. In the next sections we will apply this theorem in our proposed answering scheme for the federated adaptive data analysis situation.
@@
The first and second term in (3) shows a tradeoff in $\sigma$. The first term which is decreasing in $\sigma$, is due to over-leakage regime, where more variance for noise is required to provide better level of differential privacy. The second term which is decreasing in $\sigma$ represents under-leakage regime, where high variance of noise deteriorates data utility. In the following we derive a more manageable form for (3). First, we resolve the direct functionality of $\alpha$ in $k$, $n$ and $\sigma^{2}$. Then, we manipulate the form of this function to have $k$ as a function of $n$ and $\sigma^{2}$ for a desired given pair of $(\alpha,\beta)$. This latter form, will be directly applicable in our problem setup in the following sections.
@@
The last equation represents that $v$ is equal to Lambert $W$ function at the point $-e^{-(c+1)}$. Note that Lambert $W$ function is defined as the inverse of $y=xe^{x}$ function. Hence it satisfies: $W(z)e^{W(z)}=z$. Finally, by replacing the equivalent values and a few rearrangement, we get:
@@
For negative real numbers in $[-1/e,0)$, Lambert function has two branches, denoted by $W_{0}$ and $W_{-1}$. So we need to determine that in (6) which of them should be appeared. The branch $W_{0}$ gives values between $-1$ and 00, while $W_{-1}$ gives values in $(-\infty,-1]$. On the other hand, due to the problem constraint $10\leq\lambda<$ 1. Hence, in (6) only the usage of $W_{-1}$ is valid. Furthermore, note that variation domain of argument of $W$ in (6) lies exactly in $[-1/e,0)$ which is domain of $W_{-1}$. Finally, Substituting $\lambda^{*}$ in $f(\lambda)$, gives (5). ∎
Despite its complicated formulation in (5), $g(c)$ has a benign manner in $c$. Figure 2 depicts that for $c>10$ its growth is highly near the linear.
@@
Let $\alpha$, $\beta$ and $n$ to be given. Then $k$ is derived from $\sigma$ as follows:
@@
Letting $\sigma$ vary from 00 to greater values, we see that in (3), before a critical value of $\sigma$, the first term is selected and after that, $\alpha$ is determined by the second term. When $\alpha$ is fixed, functionality of $k$ on $\sigma$ can be derived from one of these two terms. For each $\sigma$, the term that gives lower $k$ is the correct term. ∎
In the following sections we set $\alpha=0.1$ and $\beta=0.05$, which are typical values in the literature. For this values, figure 2 shows that when $n>4\times 10^{4}$, the first term is highly near

### v2

@@ Abstract
Adaptive data analysis (ADA) involves a dynamic interaction between an analyst and a dataset owner, where the analyst submits queries sequentially, adapting them based on previous answers. This process can become adversarial, as the analyst may attempt to overfit by targeting non-generalizable patterns in the data. To counteract this, the dataset owner introduces randomization techniques, such as adding noise to the responses. This noise not only helps prevent overfitting, but also enhances data privacy. However, it must be carefully calibrated to ensure that the statistical reliability of the responses is not compromised. In this paper, we extend the ADA problem to the context of distributed datasets. Specifically, we consider a scenario where a potentially adversarial analyst interacts with multiple distributed responders through adaptive queries. We assume the responses are subject to noise, introduced by the channel connecting the responders and the analyst. We demonstrate how this noise can be opportunistically leveraged through a federated mechanism to enhance the generalizability of ADA, thereby increasing the number of query-response interactions between the analyst and the responders. We illustrate that the careful tuning of the transmission amplitude based on the theoretically achievable bounds can significantly impact the number of accurately answerable queries.
@@ I Introduction
In the era of big data, the necessity for systems that can dynamically and efficiently analyze vast and complex datasets has never been greater.One prominent approach in this domain is Adaptive Data Analysis (ADA). In ADA, the analyst aims to derive insights about an unknown distribution over a given domain by adaptively exchanging query-answer pairs with a responding mechanism that possesses a dataset of samples drawn from that distribution. However, theoretical considerations of ADA in contexts involving distributed datasets are relatively unexplored despite the growing prevalence of such datasets in modern computing environments.This paper proposes implementing ADA in distributed systems as an extension of Federated Learning (FL), which traditionally focuses on using gradient descent queries to train deep neural networks across decentralized locations. The following subsections provide a necessary examination of ADA and FL.
@@ I-A Adaptive Data Analysis
In the basic model of ADA, there is one analyst ($\mathcal{A}$) and one answering mechanism ($\mathcal{M}$), communicating with each other through error-free channels with unlimited capacities. Hence, there is no concern about practical communication limitations in this basic form of the problem. It has assumed that the answering mechanism $\mathcal{M}$ has access to the dataset $S$ with $n$ samples generated from the domain $\mathcal{X}$ by the distribution $\mathcal{D}_{\mathcal{X}^{n}}$. Albeit, in the general case, $\mathcal{D}_{\mathcal{X}^{n}}$ is not necessarily a product distribution, in the sequel, we add this assumption as it is the case in most of the basic papers in the field. So, we assume that $\mathcal{D}_{\mathcal{X}^{n}}\equiv\mathcal{P}^{n}$, where $\mathcal{P}$ is a distribution on $\mathcal{X}$. The analyst $\mathcal{A}$ does not have direct access to $S$, but can query $\mathcal{M}$ about it.Various families of queries are considered in the ADA papers, among which the most basic one is the statistical queries, studied in the pioneering work [1] and also in several subsequent works such as [2, 3]. In this paper, we also focus on this class of queries, which are formally defined below.
@@
Despite their simplicity, statistical queries encompass a substantial range of practical applications. For instance, gradient descent updates can be viewed as a vector of statistical queries. Notably, most FL papers focus on gradient descent queries due to their pivotal role in the learning process of deep neural networks. Consequently, our work can be regarded as an abstraction of these studies.
The analyst $\mathcal{A}$ can adaptively make queries based on the previous question-answer pairs. More precisely, the $i$th query, $q_{i}$, is formulated considering the sequence $(q_{1},a_{1},\dots,q_{i-1},a_{i-1})$, where $a_{j}$ is $\mathcal{M}$’s answer to the $j$th question, using the dataset $S$. The goal is to ensure that the empirical answers remain close to the true answers derived from the underlying distribution. The following definition provides the exact formulation of the accuracy criterion.
@@
In the basic model, because $\mathcal{M}$ does not know the action model of $\mathcal{A}$, it considers the worst-case scenario. Expressly, $\mathcal{M}$ assumes that $\mathcal{A}$ selects queries adversarially, aiming to overfit to the dataset $S$ and thereby violate the accuracy criterion defined by (2). On the other hand, $\mathcal{M}$ aims to answer as many queries as possible while ensuring that (2) remains satisfied. To achieve this, $\mathcal{M}$ seeks an efficient answering mechanism. In the case of non-adaptive queries (i.e., queries selected independently of observations from the data), a natural unbiased approach for $\mathcal{M}$ would be to respond using the empirical answers computed from $S$. Nevertheless, there are well-known examples demonstrating that this approach can be inefficient in adaptive scenarios. For example, see [1]. A widely adopted strategy in the research community for addressing the adaptive case is to introduce randomness into empirical answers by adding stochastic noise. The idea of adding noise comes from the data privacy framework, particularly differential privacy, which serves as a privacy-enhancing mechanism [4]. In this approach, i.i.d. noise samples from a suitable distribution are generated and added to each empirically computed answer. The variance of the noise is calibrated to ensure that the accuracy criterion in (2) is satisfied. The noise distribution is typically chosen to be either Gaussian or Laplace. In this paper, we focus on the Gaussian distribution, as we aim to leverage Gaussian channel noise as the source of randomization.
Since 2015, numerous studies have examined various aspects of ADA. In this paper, we apply one of the foundational results to our approach, rather than delving into the theoretical details. To our knowledge, no previous work has investigated the physical channel aspects of this problem. Moreover, this study is the first to explore the federated scenario within the context of ADA.
@@ I-B Federated learning
As previously noted, prior studies have not explored the physical and distributed aspects of ADA. However, the notion of distributed responders introduced in this work closely aligns with concepts in FL. To provide context, we cite selected works from the extensive FL literature that explore ideas analogous to those studied here.
In FL, a central node, called the Parameter Server (PS), interacts with multiple edge nodes, or clients. Together, these nodes work to train a shared neural network. Each client uses its local dataset to contribute to the training process. At the start of each training round, the PS broadcasts the updated weights to all nodes. Each node computes the gradient of the loss function using the current parameter values and its local dataset. Finally, the nodes send their local gradient estimations to the PS, which aggregates them to produce a more accurate global estimation.
The literature proposes various channel models and communication methods for the gradient transmission step. These approaches are generally divided into two categories: analog transmission and digital transmission. In this paper, we adopt the analog approach. With analog transmission, clients simultaneously send their estimates to the PS using uncoded transmission. This method is bandwidth-efficient because the channel is shared. Moreover, the additive property of the medium allows local estimations to naturally aggregate in the air. While this approach offers computational and storage efficiency, the aggregated signal may contain more noise and interference than digital transmission. However, these effects are not inherently detrimental. Some FL studies have utilized channel distortions to enhance client data privacy. Furthermore, [5] demonstrates that, although channel imperfections can slow the convergence rate, they may also improve the accuracy of the trained network. This suggests a nuanced relationship between data privacy and algorithm accuracy.
In the context of FL, the channel model is a critical consideration, particularly the availability of Channel State Information (CSI). Several studies, such as [6, 7, 8, 9], assume perfect global CSI at the PS and perfect local CSI at the edge devices. Some Others, including [10, 11, 12, 13, 14, 15], focus solely on perfect local CSI at the edge devices. In contrast, [16] assumes that each edge device has imperfect knowledge of its channel gain to the PS, resulting in some mismatch between the estimated and actual channel gain. Another form of imperfect CSI is examined in [15, 17], where the PS is aware only of the sum of the fading coefficients, and the edge devices lack any CSI knowledge. Finally, [5] proposes a scheme that requires no CSI at all.
@@
For a more practical approach, several works, including [13, 10, 12], consider fading effects, specifically the block-flat fading regime, where the channel condition remains constant during each communication round and may change afterward. In [6], fast fading is considered. Also [11] incorporates both path loss and small-scale fading into the model, assuming that the EPs have perfect knowledge of both types of losses.
@@ I-C Contributions and Paper Structure
In this study, we explore, for the first time, the execution of ADA over noisy communication channels. As noted earlier, the ADA framework typically introduces independent randomnesses to answers to preserve statistical validity. Thus, instead of treating channel noise as a drawback, we propose leveraging it as a natural substitute for artificially added noise.
Furthermore, we extend this approach to a distributed scenario, where an analyst—potentially adversarial—interacts with multiple EPs via a sequence of queries and responses rather than working with a single answerer. The responses are transmitted through an Additive White Gaussian Noise (AWGN) channel. We propose a response mechanism that exploits the channel noise to enhance the generalizability of ADA and optimize the number of query-response interactions.
@@ II System Model
As depicted in Figure 1, a distributed situation of ADA is considered. In this problem, we assume that each of the $L$ Edge Points (EPs) $\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{L}$ are connected to a Central Point (CP), referred to as the analyst, via a communication channel with limited capacity. Each EP, such as $\mathcal{M}_{l}$, holds a local dataset $S_{l}$ containing $n_{l}$ data samples and functions as an answering mechanism. The CP broadcasts its queries via error-free channels to all or some of them, and receives the answers from them through a wireless MAC. Note that this channel model is similar to those proposed in many recent FL studies, some of which were referenced in the previous section. This scenario can model a network of sensors communicating with a CP to infer environmental characteristics using data samples collected by the sensors.
In our proposed scheme, the answering EPs transmit analog signals with amplitudes proportional to the query answers, allowing the noisy channel to act as a stabilizing mechanism. Due to the independence of the noise realizations, each transmission experience a new randomness, aligning with the widely adopted solution for controlling bias in the basic model of the ADA problem. This approach has also been applied in some recent FL studies to preserve data privacy for the EPs.
@@ III Point-to-Point ADA without any communication constraint
In this section, we present a key theorem from [3] regarding the basic model of ADA. The theorem establishes a relationship between the variance of Gaussian noise and the accuracy parameters of the answering mechanism. While our representation differs slightly from the original, it remains mathematically equivalent. In the subsequent sections, we apply this theorem to our proposed answering scheme for the federated adaptive data analysis.
@@
The first and second terms in (3) illustrate a tradeoff in $\sigma$. The first term, which decreases with $\sigma$, is due to over-leakage regime, corresponds to the over-leakage regime, where a higher variance for noise is required to achieve a better level of differential privacy. The second term, which increases with $\sigma$, represents the under-leakage regime, where a high variance of noise deteriorates data utility. In the following, we derive a more manageable form for (3). First, we express the direct dependence of $\alpha$ on $k$, $n$ and $\sigma^{2}$. Then, we manipulate this function to express $k$ as a function of $n$ and $\sigma^{2}$ for a given pair of $(\alpha,\beta)$. This latter form, will be directly applicable in our problem setup in the following sections.
@@
The last equation shows that $v$ equals Lambert $W$ function evaluated at the point $-e^{-(c+1)}$. The Lambert $W$ function is defined as the inverse of $y=xe^{x}$ function, meaning it satisfies $W(z)e^{W(z)}=z$. Finally, by substituting the equivalent values and performing a few rearrangements, we obtain:
@@
For negative real numbers in $[-1/e,0)$, Lambert function has two branches, denoted by $W_{0}$ and $W_{-1}$. Therefore, we need to determine which terms should appear in (6). The branch $W_{0}$ gives values between $-1$ and 00, while $W_{-1}$ gives values in $(-\infty,-1]$. On the other hand, due to the problem constraint $10\leq\lambda<$ 1, only the branch of $W_{-1}$ is valid in (6). Furthermore, note that the variation domain of the argument of $W$ in (6) lies exactly in $[-1/e,0)$, which is the domain of $W_{-1}$. Finally, substituting $\lambda^{*}$ in $f(\lambda)$, gives (5). ∎
Despite its complicated formulation in (5), $g(c)$ behaves benignly with respect to $c$. Figure 2 shows that for $c>10$, its growth closely follows a linear trend.
@@
Let $\alpha$, $\beta$, and $n$ to be given. Then, $k$ is derived from $\sigma$ as follows:
@@
Letting $\sigma$ vary from 00 to larger values, we observe in (3) that, before reaching a critical value of $\sigma$, the first term dominates. After this critical value, $\alpha$ is determined by the second term. When $\alpha$ is fixed, the functional dependence of $k$ on $\sigma$ can be derived from either of these two terms. For each $\sigma$, the term that yields the smaller value of $k$ is the correct one. ∎
In the following sections, we set $\alpha=0.1$ and $\beta=0.05$, typical values in the literature. For these values, Figure 2 shows that when $n>4\times 10^{4}$, the first term closely approximates